# MIRAGE: Multimodal foundation model and benchmark for comprehensive retinal OCT image analysis

### v2

@@ Abstract
Artificial intelligence (AI) has become a fundamental tool for assisting clinicians in analyzing ophthalmic images, such as optical coherence tomography (OCT). However, developing AI models often requires extensive annotation, and existing models tend to underperform on independent, unseen data. Foundation models (FMs), large AI models trained on vast unlabeled datasets, have shown promise in overcoming these challenges. Nonetheless, available FMs for ophthalmology lack extensive validation, especially for segmentation tasks, and focus on a single imaging modality. In this context, we propose MIRAGE, a novel multimodal FM for the analysis of OCT and scanning laser ophthalmoscopy (SLO) images. Additionally, we propose a new evaluation benchmark with OCT/SLO classification and segmentation tasks. The comparison with general and specialized FMs and segmentation methods shows the superiority of MIRAGE in both types of tasks, highlighting its suitability as a basis for the development of robust AI systems for retinal OCT image analysis. Both MIRAGE and the evaluation benchmark are publicly available: https://github.com/j-morano/MIRAGE.
@@ 2.2 Ocular disease diagnosis
Table 1: Average performance of MIRAGE and other SOTA models for ocular disease diagnosis with linear probing. All models are based on the ViT-Large architecture. The best results are underlined. In each case, the performance of the best (underlined) and second best models across all datasets was compared to determine if there were statistically significant differences in average performance. To this end, the Wilcoxon signed-rank test was used. The results are presented in the table, with the significance level indicated by asterisks (*: $p<0.5$, **: $p<0.01$, ***: $p<0.001$).
@@
The results on OCT-based tasks show that MIRAGE outperformed competing models in all but one task, with statistically significant differences in 6 out of 9 cases when evaluated with the one-tailed Student’s t-test. The performance of our model was particularly high on datasets involving the diagnosis or staging of AMD and complex multi-class classification tasks. For example, on the Duke iAMD dataset (for iAMD detection), MIRAGE achieved an AUROC of 99.52%, outperforming the second-best model, DINOv2, by a significant difference of 0.39 percentage points (pp) ($p<0.01$). The same trend was observed for glaucoma staging in the small GAMMA dataset, for which MIRAGE outperformed RETFound by 5.20 pp ($p<0.01$). Smaller but still significant differences were also found for Kermany (choroidal neovascularization [CNV], drusen, and DME detection), Noor Eye Hospital (AMD and DME detection), OCTDL (6 disease detection), and OPTIMA9C (4 disease detection and AMD staging). In Harvard Glaucoma, for glaucoma detection, and OCTID (4 disease detection), MIRAGE also outperformed the other approaches, but no statistically significant differences were found. On the other hand, MIRAGE showed consistently lower performance than other SOTA approaches in OLIVES, for DME and diabetic retinopathy (DR) discrimination. These results can be attributed to the distribution of diseases in the pretraining dataset (see Table 8 in Methods). While our in-house pretraining dataset contains a high prevalence of AMD, other diseases, such as diabetes (which was highly represented in the RETFound pretraining dataset), are scarcely represented. In particular, at least 85% of the OCT scans used to train RETFound were from the Moorfields diabetic image dataset (MEH-MIDAS), while less than 2% of the scans in our dataset were from diabetic patients. Further analysis of the results (Supplementary Table 3) shows that MIRAGE performs equally well on the first scans of the patients, which are treatment-naïve. Performance declines on the later scans, which are more likely to be affected by treatment, which introduces a data shift to which our model is less robust. Notwithstanding, the overall performance of MIRAGE on OCT-based tasks was still significantly better than that of the other models, with an average AUROC of 95.59% across all OCT tasks, outperforming the second-best model, RETFound, by a significant margin of 1.15 pp ($p<0.001$). Similar differences were observed for the other metrics, with MIRAGE achieving an average AP of 92.99% (+1.64 pp, $p<0.001$) and BAcc of 84.01% (+2.51 pp, $p<0.001$).
@@ 2.4 Segmentation of retinal lesions and layers
We utilized four public and one in-house diverse datasets to validate the adaptability of our model to lesion and layer segmentation tasks in both OCT and SLO images. Specifically, we used the Duke DME [61], AROI [62], RETOUCH [63], and GOALS [64] datasets for lesion and layer segmentation in OCT, and the SGA dataset [65] (in-house) for the segmentation of geographic atrophy (GA) in SLO. For adapting the model to the segmentation tasks, we used a decoder-only fine-tuning strategy, where a ConvNeXt-based [66] decoder was trained on top of the pretrained ViT encoder [23], which is kept frozen during downstream fine-tuning. The ConvNeXt decoder was chosen based on previous work [23] and preliminary experiments (Supplementary Table 5) showing that it is well suited for medical image segmentation. This efficient strategy allows us to evaluate how well the different pretrained encoders (and thus the corresponding pretraining approaches) are suited for OCT and SLO segmentation tasks while achieving accurate segmentations. In addition to comparing our model with the DINOv2 and RETFound models previously used for classification, we also compare it with MedSAM [42], a state-of-the-art ViT-based medical image segmentation FM. All models were trained simultaneously for layer and lesion segmentation in datasets where both tasks were available (AROI and Duke DME). However, the results are reported separately for more detailed analysis.
@@
Table 2: Average performance of MIRAGE and SOTA FMs on segmentation tasks. Average and standard deviation were calculated across all different datasets. No standard deviation was calculated for SLO segmentation tasks because only one dataset was available. The performance of the two best models was compared using the Wilcoxon signed-rank test, with the resulting p-values indicated in the table by asterisks (*$p<0.05$). The best results are underlined.
@@
SLO DINOv2 61.34 $\pm$ 24.56 49.13 $\pm$ 23.88 210.02 $\pm$ 85.59
MedSAM 72.20 $\pm$ 24.05 62.22 $\pm$ 25.58 182.22 $\pm$ 86.21
MIRAGE 75.31 $\pm$ 22.16*** 65.81 $\pm$ 24.14*** 164.42 $\pm$ 95.35**
@@ 2.6 Segmentation performance comparison with specialist models
Table 3: Average performance of MIRAGE and specialist models on segmentation tasks. The standard deviation was calculated across all different datasets. No standard deviation was calculated for SLO segmentation tasks because only one dataset was available. The best results are underlined.
@@
SLO SwinUNETR-V2 74.51 $\pm$ 23.51 65.09 $\pm$ 24.88 169.03 $\pm$ 94.67
MedNeXt 77.22 $\pm$ 19.62 67.87 $\pm$ 22.19 157.15 $\pm$ 84.21
TransUNet 82.12 $\pm$ 15.92 73.69 $\pm$ 18.65 127.19 $\pm$ 57.98
nnUNet 79.41 $\pm$ 19.26 71.33 $\pm$ 21.25 136.56 $\pm$ 74.10
MIRAGE 75.31 $\pm$ 22.16 65.81 $\pm$ 24.14 164.42 $\pm$ 95.35
MIRAGE–FFT 79.36 $\pm$ 17.07 70.03 $\pm$ 19.92 136.16 $\pm$ 43.48
@@ 2.7 Qualitative analysis of segmentation results
Figure 7: Examples of segmentations from different models. The examples belong, from top to bottom, to the following datasets: Duke iAMD (cross-dataset evaluation), AROI, Duke DME, GOALS, RETOUCH, and SGA. True positives are depicted in the grayscale value of the class; false background pixels, in red; false lesion or layer pixels, in cyan; and true but wrongly classified lesion or layer pixels, in violet. The results show that our model produces precise and accurate segmentations, rarely missing or misclassifying lesions and layers. The quality of our segmentations is appreciably higher than that of nnUNet in the cross-dataset evaluation (first row), and similar in the other datasets. Compared to MedSAM and RETFound, the differences are more pronounced, as these models often misclassify pathological regions.
@@
The qualitative results are consistent with the quantitative results, showing that our model produces precise and accurate segmentations, equal to or better than nnUNet, and significantly better than the other SOTA models, here represented by RETFound, DINOv2 and MedSAM. The higher quality of the segmentations produced by MIRAGE is particularly evident in the cross-dataset setting (Figure 7, first row highlighted in green), where the models were evaluated on Duke iAMD but trained on AROI. In this cross-dataset scenario, MIRAGE demonstrates superior generalization capabilities, producing segmentations that are more accurate than those produced by the other models. In particular, nnUNet produces an incomplete segmentation of the layers, misclassifying some of the pixels (in red in Figure 7) as background, while RETFound and MedSAM, while correctly distinguishing the layers from the background in most cases, assign large areas of the image (in violet) to the wrong layer class. For the other datasets, the differences between MIRAGE and nnUNet are subtle, with both models producing generally accurate segmentations. On the other hand, RETFound shows poor overall performance, with many layers or lesions incorrectly classified as background and assigned to the wrong layer or lesion type, especially in regions with pathological signs. MedSAM, while generally more accurate than RETFound, shares some of the same problems, with the exception of the GOALS and SGA datasets, where the segmentations are nearly as accurate as those produced by MIRAGE and nnUNet.
@@ 2.8 Effect of pseudo-labeling
Table 4: Effect of pseudo-labeling. Comparison of the performance of a ViT-Base model trained on OCT alone, and OCT with pseudo-labels for the retinal layers (OCT+Layers) using the MultiMAE approach [23] on OCT classification and segmentation tasks with linear probing. The best results are underlined. In each case, the performance of the two models across all datasets for every metric was compared using the Wilcoxon signed-rank test (*: $p<0.05$, **: $p<0.01$, ***: $p<0.001$).
@@ 2.9 Effectiveness of domain-specific multimodal pretraining
Table 5: Effect of domain-specific and multimodal pretraining. Comparison of the impact of domain-specific and multimodal pretraining on the performance of a ViT-Base model in classification and segmentation tasks with linear probing. The best results are underlined. In each case, the performance of the two best models across all datasets was compared to determine if there were statistically significant differences in average performance. For this, the Wilcoxon signed-rank test was used. The results are presented in the table, with the significance level indicated by asterisks: **: $p<0.01$, ***: $p<0.001$.
@@ 2.10 Impact of model capacity
Table 6: Effect of model capacity on downstream performance. Average performance of MIRAGE based on ViT-Base or ViT-Large across all OCT and SLO classification and segmentation tasks. The best SOTA results are also shown in gray for reference: RETFound for OCT classification, SL-IN for SLO classification, and MedSAM for OCT and SLO segmentation. The best results are underlined. The Wilcoxon signed-rank test was used to compare the performance of the two models in each task, with the resulting p-values shown in the table (*: $p<0.5$, **: $p<0.01$, ***: $p<0.001$).
@@ 3 Discussion
To evaluate the generalization capabilities and the quality of the learned representations of MIRAGE, we conducted extensive evaluations on a number of clinically relevant tasks. These tasks include diagnosis and staging from OCT and SLO images and, unlike previous works [33, 27], retinal lesion and layer segmentation. For the classification of diseases and stages in OCT and SLO images, MIRAGE was compared to other SOTA FMs, including RETFound [33] and DINOv2 [21], as well as models trained using supervised learning on the ImageNet-21k dataset of natural images (SL-IN) [29]. RETFound is a unimodal FM trained with MAE [18] on a private dataset of 700k OCT images. DINOv2 was pretrained on a private dataset of 142M natural images using an SSL approach based on self-distillation. On the other hand, the ImageNet-21k dataset contains 14M images with 21k categories of natural objects. Similar to previous work on FMs [79, 27, 44, 42, 32, 80], all models were evaluated using linear probing on the target datasets. This evaluation strategy involves optimizing a single linear classifier (only about 2k parameters) on top of the frozen pretrained encoder. This allows a fair comparison of the quality of the representations learned by the models during pretraining without the need for extensive fine-tuning. The results on 11 different datasets (10 of which are public) (Figure 2 and Table 1) show that MIRAGE significantly outperforms the other models in both OCT and SLO diagnostic and staging tasks. Among the different tasks, MIRAGE was particularly effective in diagnosing and staging AMD (see, for example, the results on Duke iAMD or 9C from Figure 2), the most prevalent retinal disease worldwide (affecting 8.7% of the global population) [81]. On the other hand, the lowest relative performance of MIRAGE was observed in the discrimination of diabetes-related lesions (see the results on the OLIVES dataset from Figure 2). These results are consistent with the distribution of the diseases in the pretraining dataset, which contains a larger proportion of AMD samples ($>50\%$) compared to other diseases such as diabetes ($<2\%$). In contrast, at least 85% of the OCT scans used to train RETFound were from the Moorfields diabetic image dataset (MEH-MIDAS). Further analysis of the results (Supplementary Table 3) showed that, while MIRAGE was less robust than RETFound to the data shift introduced by patient treatment in this dataset, it achieved similar performance on treatment-naïve baseline scans of the patients, underscoring the importance of pretraining data in the generalization capabilities of models to very specialized tasks. However, an additional cross-dataset evaluation (Figure 3) shows that MIRAGE is also more robust to general dataset shifts than the other models.
@@
In recent years, the development of FMs has become a central topic in AI research in general [83, 21, 32, 44], and in medicine in particular [33, 80, 27, 84, 85]. Pivotal works such as RETFound [33] and GMAI [84] have demonstrated the potential of FMs for medical image analysis and for democratizing access to medical AI. The development of MIRAGE further extends this line of research by demonstrating the benefits of multimodal pretraining on paired medical image data for the development of more robust and generalizable FMs. Our MIRAGE model can be easily adapted to various retinal imaging tasks, including classification and segmentation in both OCT and SLO images, with minimal tuning. Through the use of efficient tuning strategies and our proposed evaluation benchmark, we showed the superior performance of MIRAGE in a wide spectrum of downstream retinal image analysis tasks. However, the applications of MIRAGE are not limited to these tasks. Similar to other FMs, MIRAGE could be effectively used as a backbone in any method that requires the extraction of high-level features from OCT or SLO images. More than a performant classification or segmentation method, MIRAGE is a general tool for retinal image analysis with a wide range of applications. In this sense, we believe that integrating domain-specific FMs like MIRAGE into AI-based systems could significantly improve their accuracy and generalizability, ultimately leading to higher-quality AI models for healthcare. This contrasts with the development of performant but non-generalizable models that are often developed, and which can lead to skepticism about the benefits of AI and limit its adoption [33]. In line with previous work [33, 42], we have made MIRAGE publicly available for research, with the goal of improving reproducibility and accelerating the progress of AI in retinal image analysis.

### v3

@@ Abstract
Artificial intelligence (AI) has become a fundamental tool for assisting clinicians in analyzing ophthalmic images, such as optical coherence tomography (OCT). However, developing AI models often requires extensive annotation, and existing models tend to underperform on independent, unseen data. Foundation models (FMs), large AI models trained on vast unlabeled datasets, have shown promise in overcoming these challenges. Nonetheless, available FMs for ophthalmology lack extensive validation, especially for segmentation tasks, and focus on a single imaging modality. In this context, we propose MIRAGE, a novel multimodal FM for the analysis of OCT and scanning laser ophthalmoscopy (SLO) images. Additionally, we propose a new evaluation benchmark with OCT/SLO classification and segmentation tasks. The comparison with general and specialized FMs and segmentation methods shows the superiority of MIRAGE in both types of tasks, highlighting its suitability as a basis for the development of robust AI systems for retinal OCT image analysis. Both MIRAGE and the evaluation benchmark are publicly available at https://github.com/j-morano/MIRAGE.
@@ 2.2 Ocular disease diagnosis
Table 1: Average performance of MIRAGE and other SOTA models for ocular disease diagnosis with linear probing. All models are based on the ViT-Large architecture. In each case, the best results (in bold) and the second best results (underlined) across all datasets were compared to determine if there were statistically significant differences in average performance. To this end, the Wilcoxon signed-rank test was used. The results are presented in the table, with the significance level indicated by asterisks (*: $p<0.5$, **: $p<0.01$, ***: $p<0.001$).
@@
The results on OCT-based tasks show that MIRAGE outperformed competing models in all but one task, with statistically significant differences in 6 out of 9 cases when evaluated with the one-tailed Student’s t-test. The performance of our model was particularly high on datasets involving the diagnosis or staging of AMD and complex multi-class classification tasks. For example, on the Duke iAMD dataset (for intermediate AMD detection), MIRAGE achieved an AUROC of 99.52%, outperforming the second-best model, DINOv2, by a significant difference of 0.39 percentage points (pp) ($p<0.01$). The same trend was observed for glaucoma staging in the small GAMMA dataset, for which MIRAGE outperformed RETFound by 5.20 pp ($p<0.01$). Smaller but still significant differences were also found for Kermany (choroidal neovascularization [CNV], drusen, and DME detection), Noor Eye Hospital (AMD and DME detection), OCTDL (6 disease detection), and OPTIMA9C (4 disease detection and AMD staging). In Harvard Glaucoma, for glaucoma detection, and OCTID (4 disease detection), MIRAGE also outperformed the other approaches, but no statistically significant differences were found. On the other hand, MIRAGE showed consistently lower performance than other SOTA approaches in OLIVES, for DME and DR discrimination. These results can be attributed to the distribution of diseases in the pretraining dataset (see Table 8 in Methods). While our in-house pretraining dataset contains a high prevalence of AMD, other diseases, such as diabetes (which was highly represented in the RETFound pretraining dataset), are scarcely represented. In particular, at least 85% of the OCT scans used to train RETFound were from the Moorfields diabetic image dataset (MEH-MIDAS), while less than 2% of the scans in our dataset were from diabetic patients. Further analysis of the results (Supplementary Table 3) shows that MIRAGE performs equally well on the first scans of the patients, which are treatment-naïve. Performance declines on the later scans, which are more likely to be affected by treatment, which introduces a data shift to which our model is less robust. Notwithstanding, the overall performance of MIRAGE on OCT-based tasks was still significantly better than that of the other models, with an average AUROC of 95.59% across all OCT tasks, outperforming the second-best model, RETFound, by a significant margin of 1.15 pp ($p<0.001$). Similar differences were observed for the other metrics, with MIRAGE achieving an average AP of 92.99% (+1.64 pp, $p<0.001$) and BAcc of 84.01% (+2.51 pp, $p<0.001$).
@@ 2.4 Segmentation of retinal lesions and layers
We utilized four public and one in-house diverse datasets to validate the adaptability of our model to lesion and layer segmentation tasks in both OCT and SLO images. Specifically, we used the Duke DME [61], AROI [62], RETOUCH [63], and GOALS [64] datasets for lesion and layer segmentation in OCT, and the SGA dataset [65] (in-house) for the segmentation of geographic atrophy (GA) in SLO. For adapting the model to the segmentation tasks, we used a decoder-only fine-tuning strategy, where a ConvNeXt-based [66] decoder was trained on top of the pretrained ViT encoder [23], which is kept frozen during downstream fine-tuning. The ConvNeXt decoder was chosen based on previous work [23] and preliminary experiments (Supplementary Table 5) showing that it is well suited for medical image segmentation. This efficient strategy allows us to evaluate how well the different pretrained encoders (and thus the corresponding pretraining approaches) are suited for OCT and SLO segmentation tasks while achieving accurate segmentations. In addition to comparing our model with the DINOv2 and RETFound models previously used for classification, we also compare it with MedSAM [42], a state-of-the-art ViT-based medical image segmentation FM. All models were trained simultaneously for layer and lesion segmentation in datasets where both tasks were available (AROI and Duke DME). However, the results are reported separately for a more detailed analysis.
@@
Table 2: Average performance of MIRAGE and SOTA FMs on segmentation tasks. Average and standard deviation were calculated across all different datasets. No standard deviation was calculated for SLO segmentation tasks because only one dataset was available. The performance of the two best models was compared using the Wilcoxon signed-rank test, with the resulting p-values indicated in the table by asterisks (*$p<0.05$). The best results are in bold, while the second best are underlined.
@@
SLO DINOv2 61.34 49.13 210.02
MedSAM 72.20 62.22 182.22
MIRAGE 75.31 65.81 164.42
@@ 2.6 Segmentation performance comparison with specialist models
Table 3: Average performance of MIRAGE and specialist models on segmentation tasks. The standard deviation was calculated across all different datasets. No standard deviation was calculated for SLO segmentation tasks because only one dataset was available. The best results are in bold, while the second best are underlined.
@@
SLO SwinUNETR-V2 74.51 65.09 169.03
MedNeXt 77.22 67.87 157.15
TransUNet 82.12 73.69 127.19
nnUNet 79.41 71.33 136.56
MIRAGE 75.31 65.81 164.42
MIRAGE–FFT 79.36 70.03 136.16
@@ 2.7 Qualitative analysis of segmentation results
Figure 7: Examples of segmentations from different models. The examples belong, from top to bottom, to the following datasets: Duke iAMD (cross-dataset evaluation, highlighted in green), AROI, Duke DME, GOALS, RETOUCH, and SGA. True positives are depicted in the grayscale value of the class; false background pixels, in red; false lesion or layer pixels, in cyan; and true but wrongly classified lesion or layer pixels, in violet. The results show that our model produces precise and accurate segmentations, rarely missing or misclassifying lesions and layers. The quality of our segmentations is appreciably higher than that of nnUNet in the cross-dataset evaluation (first row), and similar in the other datasets. Compared to MedSAM and RETFound, the differences are more pronounced, as these models often misclassify pathological regions.
@@
The qualitative results are consistent with the quantitative results, showing that our model produces precise and accurate segmentations, equal to or better than nnUNet, and significantly better than the other SOTA foundation models, here represented by RETFound, DINOv2 and MedSAM. The higher quality of the segmentations produced by MIRAGE is particularly evident in the cross-dataset setting (Figure 7, first row highlighted in green), where the models were evaluated on Duke iAMD but trained on AROI. In this cross-dataset scenario, MIRAGE demonstrates superior generalization capabilities, producing segmentations that are more accurate than those produced by the other models. In particular, nnUNet produces an incomplete segmentation of the layers, misclassifying some of the pixels (in red in Figure 7) as background, while RETFound and MedSAM, while correctly distinguishing the layers from the background in most cases, assign large areas of the image (in violet) to the wrong layer class. For the other datasets, the differences between MIRAGE and nnUNet are subtle, with both models producing generally accurate segmentations. On the other hand, RETFound shows poor overall performance, with many layers or lesions incorrectly classified as background and assigned to the wrong layer or lesion type, especially in regions with pathological signs. MedSAM, while generally more accurate than RETFound, shares some of the same problems, with the exception of the GOALS and SGA datasets, where the segmentations are nearly as accurate as those produced by MIRAGE and nnUNet.
@@ 2.8 Effect of pseudo-labeling
Table 4: Effect of pseudo-labeling. Comparison of the performance of a ViT-Base model trained on OCT alone, and OCT with pseudo-labels for the retinal layers (OCT+Layers) using the MultiMAE approach [23] on OCT classification and segmentation tasks with linear probing. The best results are in bold. In each case, the performance of the two models across all datasets for every metric was compared using the Wilcoxon signed-rank test (*: $p<0.05$, ***: $p<0.001$).
@@ 2.9 Effectiveness of domain-specific multimodal pretraining
Table 5: Effect of domain-specific and multimodal pretraining. Comparison of the impact of domain-specific and multimodal pretraining on the performance of a ViT-Base model in classification and segmentation tasks with linear probing. The best results are in bold, and the second best are underlined. In each case, the performance of the two best models across all datasets was compared to determine if there were statistically significant differences in average performance. For this, the Wilcoxon signed-rank test was used. The results are presented in the table, with the significance level indicated by asterisks (*: $p<0.05$, **: $p<0.01$, ***: $p<0.001$).
@@ 2.10 Impact of model capacity
Table 6: Effect of model capacity on downstream performance. Average performance of MIRAGE based on ViT-Base or ViT-Large across all OCT and SLO classification and segmentation tasks. The best SOTA results are also shown in gray for reference: RETFound for OCT classification, SL-IN for SLO classification, and MedSAM for OCT and SLO segmentation. The best results are in bold, and the second best are underlined. The Wilcoxon signed-rank test was used to compare the performance of the two models in each task, with the resulting p-values shown in the table (***: $p<0.001$).
@@ 3 Discussion
To evaluate the generalization capabilities and the quality of the learned representations of MIRAGE, we conducted extensive evaluations on a number of clinically relevant tasks. These tasks include diagnosis and staging from OCT and SLO images and, unlike previous works [33, 27], retinal lesion and layer segmentation. For the classification of diseases and stages in OCT and SLO images, MIRAGE was compared to other SOTA FMs, including RETFound [33] and DINOv2 [21], as well as models trained using supervised learning on the ImageNet-21k dataset of natural images (SL-IN) [29]. RETFound is a unimodal FM trained with MAE [18] on a private dataset of 700k OCT images. DINOv2 was pretrained on a private dataset of 142M natural images using an SSL approach based on self-distillation. On the other hand, the ImageNet-21k dataset contains 14M images with 21k categories of natural objects. Similar to previous work on FMs [79, 27, 44, 42, 32, 80], all models were evaluated using linear probing on the target datasets. This evaluation strategy involves optimizing a single linear classifier (only about 2k parameters) on top of the frozen pretrained encoder. This allows a fair comparison of the quality of the representations learned by the models during pretraining without the need for extensive fine-tuning. The results on 11 different datasets (10 of which are public) (Figure 2 and Table 1) show that MIRAGE significantly outperforms the other models in both OCT and SLO diagnostic and staging tasks. Among the different tasks, MIRAGE was particularly effective in diagnosing and staging AMD (see, for example, the results on Duke iAMD or 9C from Figure 2), the most prevalent retinal disease worldwide (affecting 8.7% of the global population) [81]. On the other hand, the lowest relative performance of MIRAGE was observed in the discrimination of diabetes-related lesions (see the results on the OLIVES dataset from Figure 2). These results are consistent with the distribution of the diseases in the pretraining dataset, which contains a larger proportion of AMD samples ($50\%$) compared to other diseases such as diabetes ($2\%$). In contrast, at least 85% of the OCT scans used to train RETFound were from the Moorfields diabetic image dataset (MEH-MIDAS). Further analysis of the results (Supplementary Table 3) showed that, while MIRAGE was less robust than RETFound to the data shift introduced by patient treatment in OLIVES, it achieved similar performance on treatment-naïve baseline scans of the patients, underscoring the importance of pretraining data in the generalization capabilities of models to very specialized tasks. However, an additional cross-dataset evaluation (Figure 3) shows that MIRAGE is more robust to general dataset shifts than the other models.
@@
In recent years, the development of FMs has become a central topic in AI research in general [83, 21, 32, 44], and in medicine in particular [33, 80, 27, 84, 85]. Pivotal works such as RETFound [33] and GMAI [84] have demonstrated the potential of FMs for medical image analysis and for democratizing access to medical AI. The development of MIRAGE further extends this line of research by demonstrating the benefits of multimodal pretraining on paired medical image data for the development of more robust and generalizable FMs. Our MIRAGE model can be easily adapted to various retinal imaging tasks, including classification and segmentation in both OCT and SLO images, with minimal tuning. Through the use of efficient tuning strategies and our proposed evaluation benchmark, we showed the superior performance of MIRAGE in a wide spectrum of downstream retinal image analysis tasks. However, the applications of MIRAGE are not limited to these tasks. Similar to other FMs, MIRAGE could be effectively used as a backbone in any method that requires the extraction of high-level features from OCT or SLO images. More than a performant classification or segmentation method, MIRAGE is a general tool for retinal image analysis with a wide range of applications. In this sense, we believe that integrating domain-specific FMs like MIRAGE into AI-based systems could significantly improve their accuracy and generalizability, ultimately leading to higher-quality AI models for healthcare. This contrasts with the typical development of performant but non-generalizable models, which leads to skepticism about the benefits of AI and limits its adoption [33]. In line with previous work [33, 42], we have made MIRAGE publicly available for research, with the goal of improving reproducibility and accelerating the progress of AI in retinal image analysis.